Solve by substitution of elimination 유 - 00- Х 4
Answer #1

1 5 3 -x+ y = 8 21 (1) 3 ox-y=-1 (2) 8 4 Multiplying equation (2) with 2, we get 1 3 3 1 -X-cy=-23 y = -x +2 4 2 4 Now substi